# TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu

### v1

@@
Abstract.
In the Indian subcontinent, Telugu, one of India’s six classical languages, is the most widely spoken Dravidian Language. Despite its 96 million speaker base worldwide, Telugu remains underrepresented in the global NLP and Machine Learning landscape, mainly due to lack of high-quality annotated resources. This work introduces TeSent, a comprehensive benchmark dataset for sentiment classification, a key text classification problem, in Telugu. TeSent not only provides ground truth labels for the sentences, but also supplements with provisions for evaluating explainability and fairness, two critical requirements in modern-day machine learning tasks. We scraped Telugu texts covering multiple domains from various social media platforms, news websites and web-blogs to preprocess and generate 26,150 sentences, and developed a custom-built annotation platform and a carefully crafted annotation protocol for collecting the ground truth labels along with their human-annotated rationales. We then fine-tuned several SOTA pre-trained models in two ways: with rationales, and without rationales. Further, we provide a detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models. Lastly, we curate TeEEC, Equity Evaluation Corpus in Telugu, a corpus to evaluate fairness of Telugu sentiment and emotion related NLP tasks, and provide a fairness evaluation suite for the trained classifier models. Our experimental results suggest that training with rationales may improve model accuracy, reduce bias in models, and make the explainers’ output more aligned to human reasoning.
Sentiment analysis, Benchmark, Explainable NLP, Fairness, Low-resource language, Telugu
*Major contributors, Corresponding author (Email: niladri.s@srmap.edu.in)
Author Emails: {rajkumar_vallabhaneni, ashwin_s, supriya_manna, niladri.s, vsnmshemaharshitha_cheedella, harshitha_kurakula, anandkumar_sharma, deepakraj.b, tanuj_sarkar, navaneethkrishna_bondada, shakeer_samanthapudi}@srmap.edu.in
1. Introduction
@@
Sentiment classification (Pang et al., 2002) has emerged as a prominent text classification problem due to its ubiquitous applicability. Numerous benchmark datasets for sentiment classification in high-resource languages like English (Socher et al., 2013; Maas et al., 2011; Dong et al., 2014; Wang et al., 2010; Rosenthal et al., 2017; Orbach et al., 2021) have been released to facilitate its progress. However, scarcity of quality datasets has hindered its progress in low-resource Indic languages. In fact, in their massively multilingual sentiment classification corpus, Augustyniak et al. (Augustyniak et al., 2023) included a single dataset in Indic language (Hindi) among 79 manually selected datasets of diverse languages, meeting a certain quality standard. A few sentiment corpus for Indic languages have gained popularity in recent times including those for Hindi (Akhtar et al., 2016), Bangla (Islam et al., 2023), Marathi (Kulkarni et al., 2021), multiple languages (includes Telugu) (Doddapaneni et al., 2023), Hindi-English code-mixed (Patwa et al., 2020), Malayalam-English code-mixed (Chakravarthi et al., 2020a) and Tamil-English code-mixed (Chakravarthi et al., 2020b). Like other Indic languages, Telugu sentiment corpus is rare; to the best of our knowledge, (Mukku and Mamidi, 2017) and (Marreddy et al., 2022) are the only Telugu sentiment corpus worth mentioning. We identify several concerns regarding the reliability and completeness of these corpora. Most of these corpora suffer from one or more of the following issues: (a) small sample size (fewer than 10000 instances) (Mukku and Mamidi, 2017; Chakravarthi et al., 2020a); (b) missing adequate details of annotation protocol (Marreddy et al., 2022; Kulkarni et al., 2021; Akhtar et al., 2016); (c) weak annotation, specifically less than three annotators for each sentence (Doddapaneni et al., 2023; Kulkarni et al., 2021; Patwa et al., 2020; Chakravarthi et al., 2020a; Mukku and Mamidi, 2017); (d) lack of domain diversity (Chakravarthi et al., 2020a, b; Mukku and Mamidi, 2017; Kulkarni et al., 2021; Patwa et al., 2020); and (e) reliance on machine or human translation from English (Doddapaneni et al., 2023).
@@
In the era of data-centric machine learning research, where the quality and representativeness of data take center stage, it is essential to develop resources that not only support performance but also foster explainability, inclusiveness, and linguistic richness. Achieving high accuracy is not the only target of modern-day machine learning; it expects that the models/predictions be fair and explainable. Today’s NLP applications are built over pre-trained language models, which are trained with massive open-source data with inherent societal stereotypes and biases. This often results in model outputs which favour certain group(s). Thus, ensuring fairness (Pessach and Shmueli, 2022) in model predictions has become a critical requirement. Although, several benchmarks are available for fairness evaluation in sentiment classification for English and other resource friendly languages (D´ıaz et al., 2018; Thelwall, 2018; Kiritchenko and Mohammad, 2018; Goldfarb-tarrant et al., 2023), no real initiative for the same has been taken for Telugu and other Indic languages.
@@
Explainability has become a critical pillar in machine learning research, driven by the increasing need for transparency, accountability, and human trust in AI systems—especially in high-stakes domains and under regulatory frameworks such as the GDPR’s “right to explanation” (European Parliament and Council of the European Union, ). A class of explainability methods, known as post-hoc explainers, has gained popularity for auditing model behavior with respect to predictions (Bhatt et al., 2020). However, the effectiveness of such explainers remains dependent on two key properties (Lyu et al., 2024)faithfulness, which measures how accurately an explanation reflects the model’s true reasoning process; and plausibility, which assesses how well the explanation aligns with human intuition and understanding. Sentiment classification models are increasingly being deployed in high-stakes decision making systems which have significant social and economic impact (Du et al., 2024). There is a pressing need for sentiment classification benchmarks to assess explainability aspects, and for datasets to be curated in a manner that facilitates such evaluation—an area that existing benchmarks, particularly for Telugu and other low-resource languages, have largely ignored.
@@
Keeping in mind the evolving needs of modern-day machine learning research, we introduce TeSent, the first comprehensive Telugu sentiment classification benchmark dataset which incorporates critical aspects like explainability and fairness evaluation. TeSent offers the followings.
@@
  • A labeled dataset in Telugu of 22505 high-quality examples for three-class (sentence level) sentiment classification [^2] Sentiment classification can be performed at several levels based on the granularity of the target text: aspect level, phrase level, sentence level and document level. In this work, we consider it as a sentence level task, which classifies a given sentence into a particular sentiment polarity category $\in$ positive, negative and neutral. Throughout this paper, for brevity, we write sentiment classification to mean sentence level sentiment classification unless otherwise specified., along with human annotated rationales comprising the text snippets that one considers responsible for her/his annotated label
@@
  • A detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models
We collected user comments from four major sources: YouTube, Facebook, news websites and blogs, covering approximately 20 domains, which was subjected to a preprocessing pipeline that includes removal of code-mixed information, duplicate entries, non-Telugu text, etc. We developed a custom in-house annotation software tailored to our specific design principles. We designed a robust annotation protocol where sentences were annotated by three annotators randomly selected from a gender-balanced pool of 95 undergraduate students who are native Telugu speakers and proficient in reading and understanding Telugu. The annotation process was monitored closely throughout by the set of 6 authors who are native Telugu speakers to ensure annotation accuracy. Further, post-annotation validation was conducted to ensure dataset quality. Our experimental results suggest that training models with human rationales can improve model accuracy, lead to explanations that are more closely aligned with human reasoning, and may also help reduce bias in sentiment classification models.

### v2

@@
Abstract
In the Indian subcontinent, Telugu, one of India’s six classical languages, is the most widely spoken Dravidian Language. Despite its 96 million speaker base worldwide, Telugu remains underrepresented in the global NLP and Machine Learning landscape, mainly due to lack of high-quality annotated resources. This work introduces TeSent, a comprehensive benchmark dataset for sentiment classification, a key text classification problem, in Telugu. TeSent not only provides ground truth labels for the sentences, but also supplements with provisions for evaluating explainability and fairness, two critical requirements in modern-day machine learning tasks. We scraped Telugu texts covering multiple domains from various social media platforms, news websites and web-blogs to preprocess and generate 21,119 sentences, and developed a custom-built annotation platform and a carefully crafted annotation protocol for collecting the ground truth labels along with their human-annotated rationales. We then fine-tuned several SOTA pre-trained models in two ways: with rationales, and without rationales. Further, we provide a detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models. Lastly, we curate TeEEC, Equity Evaluation Corpus in Telugu, a corpus to evaluate fairness of Telugu sentiment and emotion related NLP tasks, and provide a fairness evaluation suite for the trained classifier models. Our experimental results suggest that training with human rationales improves model accuracy and models’ alignment with human reasoning, but does not necessarily reduce bias.
TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu
Vallabhaneni Raj Kumar1, Ashwin S1, Supriya Manna3,‡, Niladri Sett1,*, Cheedella V S N M S Hema Harshitha1, Kurakula Harshitha1, Anand Kumar Sharma1, Basina Deepakraj1, Tanuj Sarkar1, Bondada Navaneeth Krishna1, Samanthapudi Shakeer1
footnotetext: 1 SRM University AP.  2 Techmojo Solutions Private Ltd.  3 University of Maryland, College Park, USA.  4 Razorpay, India.  5 PayPal, India.  6 DevRev, India.  7 Kelp Global.  * Corresponding author.  Most work was done while SM was at SRM University AP.
1 Introduction
@@
Sentiment classification Pang et al. (2002) has emerged as a prominent text classification problem due to its ubiquitous applicability. Numerous benchmark datasets for sentiment classification in high-resource languages like English Socher et al. (2013); Maas et al. (2011); Dong et al. (2014); Wang et al. (2010); Rosenthal et al. (2017); Orbach et al. (2021) have been released to facilitate its progress. However, scarcity of quality datasets has hindered its progress in low-resource Indic languages. In fact, in their massively multilingual sentiment classification corpus, Augustyniak et al. Augustyniak et al. (2023) included a single dataset in Indic language (Hindi) among 79 manually selected datasets of diverse languages, meeting a certain quality standard. A few sentiment corpus for Indic languages have gained popularity in recent times including those for Hindi Akhtar et al. (2016), Bangla Islam et al. (2023), Marathi Kulkarni et al. (2021), multiple languages (includes Telugu) Doddapaneni et al. (2023), Hindi-English code-mixed Patwa et al. (2020), Malayalam-English code-mixed Chakravarthi et al. (2020a) and Tamil-English code-mixed Chakravarthi et al. (2020b). Like other Indic languages, Telugu sentiment corpus is rare; to the best of our knowledge, Mukku and Mamidi (2017) and Marreddy et al. (2022) are the only Telugu sentiment corpus worth mentioning. We identify several concerns regarding the reliability and completeness of these corpora. Most of these corpora suffer from one or more of the following issues: (a) small sample size (fewer than 10000 instances) Mukku and Mamidi (2017); Chakravarthi et al. (2020a); (b) missing adequate details of annotation protocol Marreddy et al. (2022); Kulkarni et al. (2021); Akhtar et al. (2016); (c) weak annotation, specifically less than three annotators for each sentence Doddapaneni et al. (2023); Kulkarni et al. (2021); Patwa et al. (2020); Chakravarthi et al. (2020a); Mukku and Mamidi (2017); (d) lack of domain diversity Chakravarthi et al. (2020a, b); Mukku and Mamidi (2017); Kulkarni et al. (2021); Patwa et al. (2020); and (e) reliance on machine or human translation from English Doddapaneni et al. (2023).
@@
In the era of data-centric machine learning research, where the quality and representativeness of data take center stage, it is essential to develop resources that not only support performance but also foster explainability, inclusiveness, and linguistic richness. Achieving high accuracy is not the only target of modern-day machine learning; it expects that the models/predictions be fair and explainable. Today’s NLP applications are built over pre-trained language models, which are trained with massive open-source data with inherent societal stereotypes and biases. This often results in model outputs that favour certain group(s). Thus, ensuring fairness Pessach and Shmueli (2022) in model predictions has become a critical requirement. Although, several benchmarks are available for fairness evaluation in sentiment classification for English and other resource friendly languages Díaz et al. (2018); Thelwall (2018); Kiritchenko and Mohammad (2018); Goldfarb-tarrant et al. (2023), no real initiative for the same has been taken for Telugu and other Indic languages.
@@
Explainability has become a critical pillar in machine learning research, driven by the increasing need for transparency, accountability, and human trust in AI systems—especially in high-stakes domains and under regulatory frameworks such as the GDPR’s “right to explanation” European Parliament and Council of the European Union (2016). A class of explainability methods, known as post-hoc explainers, has gained popularity for auditing model behavior with respect to predictions Bhatt et al. (2020). However, the effectiveness of such explainers remains dependent on two key properties Lyu et al. (2024)faithfulness, which measures how accurately an explanation reflects the model’s true reasoning process; and plausibility, which assesses how well the explanation aligns with human intuition and understanding. Sentiment classification models are increasingly being deployed in high-stakes decision making systems which have significant social and economic impact Du et al. (2024). There is a pressing need for sentiment classification benchmarks to assess explainability aspects, and for datasets to be curated in a manner that facilitates such evaluation—an area that existing benchmarks, particularly for Telugu and other low-resource languages, have largely ignored.
@@
Keeping in mind the evolving needs of modern-day machine learning research, we introduce TeSent, the first comprehensive Telugu sentiment classification benchmark dataset which incorporates critical aspects like explainability and fairness evaluation. TeSent offers the following.
@@
  • A labeled dataset in Telugu of 21,119 high-quality examples for three-class (sentence level) sentiment classification [^2] Sentiment classification can be performed at several levels based on the granularity of the target text: aspect level, phrase level, sentence level and document level. In this work, we consider it as a sentence-level task, which classifies a given sentence into a particular sentiment polarity category $\in$ positive, negative and neutral. Throughout this paper, for brevity, we write sentiment classification to mean sentence level sentiment classification unless otherwise specified., along with human annotated rationales comprising the text snippets that one considers responsible for her/his annotated label
@@
  • A detailed plausibility and faithfulness evaluation suite, which exploits the rationales, for six widely used post-hoc explainers applied on the trained models to investigate models’ alignment with human understanding from a human-understandable perspective when trained with rationales.
We collected user comments from four major sources: YouTube, Facebook, news websites, and blogs, covering approximately 20 domains, which was subjected to a preprocessing pipeline that includes removal of code-mixed information, duplicate entries, non-Telugu text, etc. We developed a custom in-house annotation software tailored to our specific design principles. We designed a robust annotation protocol where sentences were annotated by three annotators randomly selected from a gender-balanced pool of 95 undergraduate students who are native Telugu speakers and proficient in reading and understanding Telugu. The annotation process was monitored closely throughout by the set of 6 authors who are native Telugu speakers to ensure annotation accuracy. Further, post-annotation validation was conducted to ensure dataset quality. Our experimental results suggest that training models with human rationales can improve model alignment with human understanding, but may not necessarily improve fairness.